4,13,38........ Next term
(A) 58 (B)96. (C) 111 (D) 123














Answer by rothauserc(4718)

We are working with the sum of the squares of prime numbers
:
partial list of primes is 2, 3, 5, 7, 11, ...
:
x(n) = sum of p(i)^2, where p is a prime, i = 1, n
:
x(4) = 4 + 9 + 25 + 49 = 87
:
I do not see 87 as a solution
: